# Exploring non-Euclidean photonics: Pseudosphere microlaser

### v1

@@ V Solutions of the wave equation with Dirichlet boundary conditions
Following the ”effective index approximation” for curved layers proposed in Ref. [9], we assume the light is well-confined within the layer and propagates with an effective index $n_{eff}$, which accounts for the varying refractive indices in the direction perpendicular to the surface. Furthermore the numerical simulations in Sec. VI suggest that the tractoid exhibits two distinct classes of modes: Transverse Electric modes or TE-modes (resp. Transverse Magnetic modes or TM-modes) for which the electric (resp. magnetic) is parallel to the tractoid surface. Experimental results also indicate that the lasing modes of the tractoid microlaser are TE-modes (with the electric field in the plane $Z=0$). Hence we can resort to the 2D scalar Helmholtz equation
@@
where $\Delta_{s}$ is the Laplacian operator restricted to the tractoid surface, and the wavefunction $\psi$ corresponds to the component of the magnetic (resp. electric) field normal to the surface for TE-modes (resp. TM-modes).
@@
In Sec. V.1, we separate variables in order to write Eq. (18) as a Schrödinger-like equation along the radial coordinate. The localization of the wave at the boundary $u=0$ can then be interpreted thanks to an effective potential. Then in Sec. V.2, we derive the analytical solutions of Eq. (18), which are compared to the analytical solutions of the flat disk in Sec. V.3. As the analytical solutions of the tractoid are not easy to handle, we propose an efficient numerical method in Sec. V.4 to provide a large number of eigenvalues. Finally we show that their distribution is consistent with an integrable system.
@@ V.1 Schrödinger-like equation
This equation can be solved by variable separation using the ansatz $\psi(x,y)=\zeta(y)e^{imx}$ with $m\in\mathbb{Z}$ due to the periodic conditions in the $x$-direction. By inserting the expression of $\psi$ in the wave equation (21), we obtain the equation satisfied by $\zeta$
@@
where $\mu^{2}=(kR_{0})^{2}-\frac{1}{4}$. This equation corresponds to a Schrödinger equation for a particle trapped in the potential well $V_{eff}^{(m)}(\eta)=m^{2}e^{2\eta}$ plotted in Fig. 7. Furthermore it can be transformed into a Bessel equation with imaginary index, as shown in the next paragraph.
Figure 7: Effective potential well emerging from the Schrödinger-like equation (24) for the truncated tractoid (black solid line). The red line is the wavefunction corresponding to $m=60$ and $p=3$. $\eta=0$ corresponds to the boundary $u=0$.
@@ V.2 Analytic solutions
$\frac{d^{2}\gamma}{dw^{2}}+\frac{1}{w}\frac{d\gamma}{dw}+\left[-1+\frac{(kR_{0})^{2}-\frac{1}{4}}{w^{2}}\right]\gamma(w)=0$ (25)
which is actually a modified Bessel equation for the complex parameter $i\mu$ with $\mu^{2}=(kR_{0})^{2}-1/4$. This equation has two linearly independent solutions which are called the modified Bessel function of the first and the second kind $K_{i\mu}$ and $I_{i\mu}$. They are related to the Bessel functions of the first and the second kind by
@@
with $w$ real. A general solution of the wave equation can thus be written as a linear combination of $I_{i\mu}$ and $K_{i\mu}$. However, $I_{i\mu}$ diverges for $w\to\infty$ and $I_{i\mu}$ diverges for $w\to 0$. Therefore, if we consider the whole tractoid with $u\to\infty$ (ie. $w\to\infty)$, only the $K_{i\mu}$ solution should remain. Consequently, the full wavefunction is given by
@@
with $C\in\mathbb{C}$ a constant. For simplicity, we consider Dirichlet boundary condition at $u=0$, which means
@@
For a given integer $m$, the real $k$ values satisfying Eq. (29) are denoted as $k_{m,p}$ with $p=1,2,\ldots$ the radial order of the root. Finding these roots $k_{m,p}$ is still an open problem in spite of the broad bibliography. Recently, it became possible to obtain an analytical solution of this equation for large values of $k$ [17]. Nevertheless, the matricial method developed in Sec. V.4 provides an efficient way to approximate these roots. Indeed, the eigenvalues computed in Sec. V.4 are compared to the roots of Eq. (29) around $m=130$ computed by the FindRoot command on Mathematica and plotted in Fig. 8. The agreement between both spectra is excellent, with a relative difference smaller than $4\times 10^{-6}$ for the first branch $p=1$ and around $2\times 10^{-5}$ for the third branch $p=3$.
Figure 8: Comparison between the analytical spectrum of the tractoid computed with Mathematica (open squares) and the numerical spectrum computed using the matricial approach developped in Sec. V.4 (dots).
@@ V.3 Comparison with the flat disk
Solutions of the wave equation for a disk of radius $R_{0}$ with Dirichlet boundary conditions are the Bessel function of the first kind $J_{m}$ with $m\in\mathbb{N}$. The spectrum is real positive and is given by the boundary conditions
@@
A major difference with Eq. (29) for the tractoid is that the wavenumber $k$ now appears in the order $i\mu$ of the modified Bessel function and not in the argument as it was the case for the disk. The spectra of the disk and the tractoid are plotted in figure 9. When $m$ is sufficiently large, and $p$ small, modes of the disk and the tractoid are very close to each other. For $p=1$ and $m\simeq 100$, the relative difference between modes of the disk and the tractoid is around 0.2 %. When $p$ gets larger, the difference also increases. Actually for a large azimuthal number $m$ and a small radial number $p$, the wave is mainly located along the circular boundary of the tractoid [cf. Fig. 10(a)] where the tractoid surface looks similar to a flat disk.
V.4 Numerical resolution of the wave equation
@@
Since the roots of Eq. (29) are not easy to compute, in this last section, we propose a convenient numerical method to solve the Schrödinger-like Eq. (24) for the effective potential of Fig. 7 corresponding to the tractoid. This method is also applicable to many other potentials.
We start by spatially discretizing the problem. We consider $N+2$ equally spaced points $0=\eta_{0}<\eta_{1}<...<\eta_{N}<\eta_{N+1}=\eta_{max}$ and we note $\xi_{i}=\xi(\eta_{i})$ for $i\in[\![0,N+1]\!]$. For the results presented in the following, $N$ is equal to 9400.
@@
For comparison with previous paragraphs, we consider Dirichlet boundary conditions at $\eta=0$ (corresponding to $u=0$). As discussed before, the wavefunctions are expected to be localized near the potential minimum at $\eta=0$, and not sensitive to the boundary condition at $\eta=\eta_{max}$ (corresponding to a tractoid truncated at $u=u_{max}$). Therefore we also apply Dirichlet boundary condition at $\eta=\eta_{max}$ for the sake of simplicity. To summarize, we have to solve Eq. (24) with the following Dirichlet boundary conditions
$\xi(\eta=0)=\xi(\eta_{max})=0$ (31)
where $e^{\eta_{max}}=\cosh(u_{max})$.
For a sufficiently small spatial step $h=\frac{\eta_{N+1}-\eta_{0}}{N+1}$, the second order derivative is well approximated by the discretization
@@
for $i\in[\![1,N]\!]$. Therefore, we obtain a system of linear equations that can be written with matrix notations as
$(T+V)\,\vec{\xi}=H\vec{\xi}=\mu^{2}\vec{\xi}$ (33)
@@
where the notation $H=T+V$ has been used. $T$ is a tridiagonal matrix which accounts for the kinetic energy term described by expression (32) and the effective potential is represented by the diagonal matrix $V$. The explicit expressions of the matrices are given in App. E. The operator $H$ is then equivalent to the Hamiltonian operator of a quantum mechanical system.
@@
Figure 10: (a) Differents views of the $(10,4)$-mode (top) and the $(111,2)$-mode (bottom) calculated by the matricial approach of Sec. V.4. The wavefunctions are real-valued. (b) Nearest Neighbour Spacing Distribution calculated based on the 4000 modes found between $kR_{0}=75$ and $kR_{0}=150$. The Wigner distribution from the Gaussian Orthogonal Ensemble (GOE) is plotted with a dashed black line, and the Poissonian distribution with a solid red line.
@@
The problem of finding the wave functions and the corresponding wavenumbers $kR_{0}$ has been turned into finding the eigenvectors and the eigenvalues of the tridiagonal matrix $H$. It can be done by using the Relatively Robust Representations method implemented on Python in the scipy library. Similarly to the disk, modes are denominated by two integers, $m$ and $p$. For a given azimuthal number $m\in\mathbb{Z}$, there are several solutions indexed by the radial number $p\in\mathbb{N}^{*}$. Some examples of eigen wavefunctions are shown in Fig. 10(a). They look like WGM in a flat disk.
@@
Using this numerical approach for $u_{max}=3$, we found about 4000 eigenenergies from $kR_{0}=75$ to $kR_{0}=150$. By comparison with the Weyl formula (see App. F), we then obtained 98.5% of the eigenvalues in this range. The Nearest Neighbour Spacing Distribution (NNSD) was calculated based on the method described in App. G and is plotted in Fig. 10(b). It exhibits a Poissonian distribution as expected for integrable systems [18].
@@
To conclude this section, as the tractoid is rotationally symmetric, its wave equation is separable. Its radial part verifies a modified Bessel equation of imaginary order. Its solutions are similar to the wavefunctions of the flat disk, at least near the bottom of the effective potential, at the boundary $u=0$. This corroborates and enlightens the observation of lasing WGMs in experiments.
@@ VI Mode computation in the dielectric tractoid cavity
Due to their rotational symmetry, the modes of the disk and the tractoid are two-fold degenerate. Therefore we restric the simulation domain to half a tractoid (and half a disk) and to even wavefunctions, which corresponds to impose Neumann boundary conditions along the plane $X=0$. The source is a linear dipole positioned in the plane $X=0$ at $u=0.23$, which is very close to the boundary. In Figs. 11 and 12, the source is oriented parallel to the $Y$ axis, and thus excites preferentially TE-polarized modes. However if the source is parallel to the $Z$ axis, the simulation does not yield any useable signal.
@@
Fig. 12 presents the four modes inferred from the tractoid simulations with a source at a central frequency Re$(kR_{0})\simeq 100$. The profiles of these wavefunctions are in complete agreement with the results of Sec. V. Each mode belongs to a different branch of the spectrum of Fig. 11. The modes of branches 1, 2, and 4 are TE-polarized with $p=1$, 2, and 3 radial excitations respectively, while branch 3 is TM-polarized with a single radial excitation $p=1$. For the disk, only TE-modes are excited with such a source.
@@ VII Conclusion
In summary, the tractoid is integrable thanks to it rotational symmetry and in spite of the exponential divergence of ray trajectories induced by the negative curvature. When breaking the rotational symmetry, however, chaotic dynamics is expected, which will be the focus of a further exploration in non-Euclidean photonics.
@@ Acknowledgements.
This work was done with the C2N micro nanotechnologies platform. It was supported by the French RENATECH network, by the departement de l’Essonne, and by the French state funds managed by the ANR within the Investissements d’Avenir programme under reference ANR-19-CE09-0017 FRONTAL. S.B. acknowledges support for the Chair in Photonics from Région Grand-Est, Département de la Moselle, European Regional Development Fund, CentraleSupélec, Fondation CentraleSupélec, and the Eurometropole de Metz.
@@
J.-F. Audibert is acknowledged for the fabrication of a centimeter-scale tractoid with a 3D printer, Y.-H. Melka and O. Ivashtenko for preliminary works on spherical squares, and A. Comtet for Sec. V.2.
Appendix A Conformal mappings of the tractoid
In this Appendix, we describe two conformal mappings of the tractoid, different from the Poincaré half-plane, which are not used in this paper but which may be useful depending on the context.
A.1 Poincaré Disk
The Poincaré disk is another well-known representation of hyperbolic geometry [4]. It corresponds to the unit disk $\mathbb{D}=\{(x_{1},x_{2})\in\mathbb{R}^{2},x_{1}^{2}+x_{2}^{2}=1\}$ endowed with the metric
$ds_{disk}^{2}=4R_{0}^{2}\frac{dx_{1}^{2}+dx_{2}^{2}}{1-x_{1}^{2}-x_{2}^{2}}$ (34)
There is a one-to-one mapping of the Poincaré half-plane onto the Poincaré disk which is given in complex coordinates by
$w=\frac{-z+i}{z+i}$ (35)
where $z=x+iy$ (resp. $w=x_{1}+ix_{2}$) the complex coordinate in the Poincaré half-plane (resp. Poincaré disk). Like in the Poincaré half-plane, geodesics are simple in the Poincaré disk. There are only two kinds of geodesics: the diameter and arc cercle perpendicular to the boundary of the disk (cf. Fig. 13). In fact, the diameter can also be seen has an arc circle of infinite radius which is perpendicular to the boundary of the disk. Finally, it should be noted that, similarly to the Poincaré half-plane, some parts of the Poincaré disk cannot be mapped onto the tractoid. Actually, only the disk of center $\left(-\frac{1}{2},0\right)$ and radius $\frac{1}{2}$ can be mapped onto the tractoid (cf. Fig. 13).
Figure 13: Schematic of the Poincaré disk. The hatched zone is the subspace of the disk which cannot be mapped onto the tractoid. The green and brown lines correspond to the two kinds of geodesics.
A.2 Conformal mapping for surfaces of revolution
Another useful conformal mapping of the tractoid has been studied in Ref. [13]. This transformation works for any surface of revolution. It is given by
[…]
where $A\in\mathbb{R}$ is a constant and the $\pm$ sign corresponds to two different transformations. From now on, we take $A=eR_{0}$ with $e$ the Euler’s number. The induced metric is
$ds^{\prime 2}$ $=n^{2}(r)(dr^{2}+r^{2}d\varphi^{2})$ (37)
$n(r)$ $=\frac{\sqrt{G(u(r))}}{r}$ (38)
A physical interpretation of this transformation is that the surface of revolution is mapped into a 2D flat plane with a spatially-varying optical index $n(r)$. In particular, the rotational symmetry is preserved under this transformation which will be convenient to derive the wave equation by variable separation. Applying this method to the specific case of the tractoid, the conformal mapping is defined for all $u\in[0,+\infty)$ by
$r(u)=R_{0}e^{1-\cosh(u)}$ (39)
where the -- sign has been chosen in the exponential term. Then, the tractoid is mapped onto a disk of radius $R_{0}$. The point $u=+\infty$ is mapped to the center of the disk and the boundary $u=0$ becomes the boundary of the disk. The varying optical index is therefore defined for all $r\in(0,R_{0}]$ by
$n(r)=\frac{R_{0}}{r\left[1-\ln\left(\frac{r}{R_{0}}\right)\right]}$ (40)
The optical index is not defined at the center of the disk because of the singular behavior of the tractoid at $u=+\infty$. In practice, the microlasers are truncated at $u=u_{max}$. Then, the truncated tractoid transforms as a ring with inner radius $r_{min}=R_{0}e^{1-\cosh(u_{max})}$ and outer radius $r_{max}=R_{0}$. The index profile of the mapped ring is shown in Fig. 14.
Figure 14: Radial optical index profile of the mapped tractoid, following Eq. (40). Inset: 2D spatial distribution of the optical index.

### v3

@@ V Solutions of the wave equation with Dirichlet boundary conditions
Following the ”effective index approximation” for curved layers proposed in Ref. [13], we assume that the light is well-confined within the layer and propagates with an effective index $n_{eff}$, which accounts for the varying refractive indices in the direction perpendicular to the surface. Furthermore Ref. [24] demonstrated that the tractoid surface should support two distinct polarization classes of modes, Transverse Electric modes or TE-modes (resp. Transverse Magnetic modes or TM-modes) for which the electric (resp. magnetic) is parallel to the tractoid surface. Actually all the necessary assumptions of Ref. [24] are valid for the tractoid, namely it is a surface of revolution, its main curvature $H$ is null, and the lasing wavelength is much smaller than the scale of the curvature. Experimental results also indicate that the lasing modes of the tractoid microlaser are TE-modes (with the electric field in the plane $Z=0$). Hence we can resort to the 2D scalar Helmholtz equation
@@
where $\Delta_{s}$ is the Laplace operator restricted to the tractoid surface, and the wave function $\psi$ corresponds to the component of the magnetic (resp. electric) field normal to the surface for TE-modes (resp. TM-modes).
@@
In Sec. V.1, we separate variables in order to write Eq. (18) as a Schrödinger-like equation along the radial coordinate. The localization of the wave at the boundary $u=0$ can then be interpreted through an effective potential. Then in Sec. V.2, we derive the analytical solutions of Eq. (18), which are compared to the analytical solutions of the flat disk in Sec. V.3. As the analytical solutions of the tractoid are not easy to handle, we propose an efficient numerical method in Sec. V.4 to provide a large number of eigenvalues. Finally we show that their distribution is consistent with an integrable system.
@@ V.1 Schrödinger-like equation
This equation can be solved by variable separation using the ansatz $\psi(x,y)=\zeta(y)e^{imx}$ with $m\in\mathbb{Z}$. By inserting the expression of $\psi$ in the wave equation (21), we obtain the equation satisfied by $\zeta$
@@
where $\mu^{2}=(kR_{0})^{2}-\frac{1}{4}$. This equation corresponds to a Schrödinger equation for a particle trapped in the potential well $V_{eff}^{(m)}(\eta)=m^{2}e^{2\eta}$. It is plotted for $m=60$ in Fig. 7 together with the occupation probability density $\psi(u,v)^{2}$ with $(m,p)=(60,3)$, which is similar to that for the corresponding flat disk near the bottom of the effective potential, at the boundary $u=0$, ie. $\eta=0$.
Furthermore Eq. (24) can be transformed into a Bessel equation with imaginary index, as shown in the next paragraph.
Figure 7: Effective potential well, $V_{eff}^{(m)}(\eta)=m^{2}e^{2\eta}$ for $m=60$ emerging from the Schrödinger-like equation (24) for the truncated tractoid (black solid line). $\eta=0$ corresponds to the boundary $u=0$. The red line is the real-valued wave function corresponding to $m=60$ and $p=3$. With a $\mu^{2}$ offset, it oscillates around the horizontal dotted line which is positioned at the eigen-energy $\mu^{2}$.
@@ V.2 Analytic solutions
$\frac{d^{2}\gamma(\omega)}{dw^{2}}+\frac{1}{w}\frac{d\gamma(\omega)}{dw}-\left[1+\frac{\frac{1}{4}-(kR_{0})^{2}}{\omega^{2}}\right]\gamma(w)=0$ (25)
which is actually the equation for a modified Bessel function with index $\tilde{\mu}$, where $\tilde{\mu}^{2}=1/4-(kR_{0})^{2}$ and $\tilde{\mu}=i\sqrt{(kR_{0})^{2}-1/4}\equiv i\mu$, which is purely imaginary for $kR_{0}>1/4$. This equation has two linearly independent solutions which are called the modified Bessel function of the first and the second kind $K_{i\mu}$ and $I_{i\mu}$. They are related to the Bessel functions of the first and the second kind by
@@
A general solution of the wave equation can thus be written as a linear combination of $I_{i\mu}$ and $K_{i\mu}$. However, $I_{i\mu}(w)$ diverges for $w\to\infty$, while $K_{i\mu}(w)$ is finite. Therefore, if we consider the whole tractoid with $u\to\infty$ (ie. $w\to\infty)$, only the $K_{i\mu}$ solution should remain:
@@
with $C\in\mathbb{C}$ a constant. For simplicity, we consider the Dirichlet boundary condition at $u=0$, that is
@@
For a given integer $m$, the real $k$ values satisfying Eq. (29) are denoted as $k_{m,p}$ with $p=1,2,\ldots$ the radial order of the root. Finding these roots $k_{m,p}$ is still an open problem in spite of the broad bibliography. Recently, it became possible to obtain an analytical solution of this equation for large values of $k$ [25]. Nevertheless, the matrix method developed in Sec. V.4 provides an efficient way to approximate these roots. Indeed, the eigenvalues computed in Sec. V.4 are compared to the roots of Eq. (29) around $m=130$ computed by the FindRoot command on Mathematica and plotted in Fig. 8. The agreement between both spectra is excellent, with a relative difference smaller than $4\times 10^{-6}$ for the first branch $p=1$ and around $2\times 10^{-5}$ for the third branch $p=3$.
Figure 8: Comparison between the analytical spectrum of the tractoid computed with Mathematica (open squares) and the numerical spectrum computed using the matrix approach developed in Sec. V.4 (dots).
@@ V.3 Comparison with the flat disk
Solutions of the radial wave equation for a disk of radius $R_{0}$ with Dirichlet boundary conditions are the Bessel function of the first kind, that is, $\psi(r,\varphi)=J_{m}(kr)\phi(m\varphi)$, with $\phi(x)=\cos x$ or $\phi(x)=\sin x$ in polar coordinates $(r,\varphi)$ with $m\in\mathbb{N}$. The spectrum is real positive and is given by the boundary conditions
@@
A major difference with Eq. (29) for the tractoid is that there the wavenumber $k$ appears in the index $i\mu$ of the modified Bessel function and not in the argument, as in Eq. (30). The spectra of the disk and the tractoid are plotted in Fig. 9. When $m$ is sufficiently large, and $p$ small, modes of the disk and the tractoid are very close to each other. For $p=1$ and $m\simeq 100$, the relative difference between modes of the disk and the tractoid is around 0.2 %. When $p$ gets larger, the difference also increases. Actually for a large azimuthal number $m$ and a small radial number $p$, the wave is mainly located along the circular boundary of the tractoid [cf. Fig. 10(a)] where the tractoid surface looks similar to a flat disk.
V.4 Numerical procedure to solve the wave equation
@@
Since the roots of Eq. (29) are not easy to compute, we employed a convenient numerical finite-element method to solve the Schrödinger-like Eq. (24) for the effective potential of Fig. 7 corresponding to the tractoid, which is explained in App. E.
@@
For comparison with previous paragraphs, we consider Dirichlet boundary conditions at $\eta=0$ (corresponding to $u=0$). As discussed before, the wave functions are expected to be localized in the potential well at $\eta=0$, and to be insensitive to the boundary condition at $\eta=\eta_{max}$ (corresponding to a tractoid truncated at $u=u_{max}$). Therefore to simplify, we also apply Dirichlet boundary condition at $\eta=\eta_{max}$. To summarize, we have to solve Eq. (24) with the Dirichlet boundary conditions $\xi(\eta=0)=\xi(\eta_{max})=0$, where $e^{\eta_{max}}=\cosh(u_{max})$.
@@
After a discretization protocol explained in App. E, we obtain a system of linear equations that can be written with matrix notations as
$(\mathcal{T}+\mathcal{V})\,\vec{\xi}=\mathcal{H}\vec{\xi}=\mu^{2}\vec{\xi}$ (31)
@@
where the notation $\mathcal{H}=\mathcal{T}+\mathcal{V}$ has been used. $\mathcal{T}$ is a tridiagonal matrix which accounts for the kinetic energy term described by expression (62) and the effective potential is represented by the diagonal matrix $\mathcal{V}$. The operator $\mathcal{H}$ is then equivalent to the Hamiltonian operator of a quantum mechanical system.
@@
Figure 10: (a) Different views of the $(10,4)$-mode (top) and the $(111,2)$-mode (bottom) calculated by the matrix approach of Sec. V.4. The wave functions are real-valued. (b) Nearest-Neighbour Spacing Distribution calculated based on the 4000 modes found between $kR_{0}=75$ and $kR_{0}=150$. The Wigner surmise for random matrices from the Gaussian Orthogonal Ensemble (GOE) is plotted with a dashed black line, and the Poissonian distribution with a solid red line.
@@
The problem of finding the wave functions and the corresponding wavenumbers $kR_{0}$ has been turned into finding the eigenvectors and the eigenvalues of the tridiagonal matrix $\mathcal{H}$. It can be done by using the Relatively Robust Representations method implemented on Python in the scipy library. Similarly to the disk, modes are denominated by two integers, $m$ and $p$. For a given azimuthal number $m\in\mathbb{Z}$, there are several solutions indexed by the radial number $p\in\mathbb{N}^{*}$. Some examples of wave functions are shown in Fig. 10(a). They ressemble WGMs in a flat disk.
@@
Using this numerical approach for $u_{max}=3$ and $N=9400$ discretization points, we found about 4000 eigenenergies from $kR_{0}=75$ to $kR_{0}=150$. By comparison with the Weyl formula (see App. F), we then obtained 98.5% of the eigenvalues in this range. The Nearest-Neighbour Spacing Distribution (NNSD) was calculated based on the method described in App. G and is plotted in Fig. 10(b). It exhibits a Poissonian distribution as expected for integrable systems [26].
@@
To conclude this section, as the tractoid is rotationally symmetric, its wave equation is separable. Its radial part solves a modified Bessel equation. Its solutions are similar to the wave functions of the flat disk, at least near the bottom of the effective potential, at the boundary $u=0$. This corroborates and enlightens the observation of lasing WGMs in experiments.
@@ VI Mode computation in the dielectric tractoid cavity
Due to their rotational symmetry, the modes of the disk and the tractoid are two-fold degenerate. Therefore we restrict the simulation domain to half a tractoid (and half a disk) and to even wave functions, which corresponds to imposing Neumann boundary conditions along the plane $X=0$. The source is a linear dipole positioned in the plane $X=0$ at $u=0.23$, which is very close to the boundary. In Figs. 11 and 12, the source is oriented parallel to the $Y$ axis, and thus excites preferentially TE-polarized modes. However if the source is parallel to the $Z$ axis, the simulation does not yield any useable signal.
@@
Fig. 12 presents the four modes inferred from the tractoid simulations with a source at a central frequency Re$(kR_{0})\simeq 100$. The profiles of these wave functions are in complete agreement with the results of Sec. V. Each mode belongs to a different branch of the spectrum of Fig. 11. The modes of branches 1, 2, and 4 are TE-polarized with $p=1$, 2, and 3 radial excitations respectively, while branch 3 is TM-polarized with a single radial excitation $p=1$. This splitting of polarizations is expected since the tractoid surface fulfills the corresponding assumptions of the model in Ref. [24]. For the disk, only TE-modes are excited with such a source.
@@
Finally, Fig. 13 shows that this 3D numerical spectrum is very close to the analytic spectrum obtained in Sec. V.2 and V.4, which validates a posteriori the use of Dirichlet boundary conditions in the analytic treatment, consistent high quality factor WGMs [27].
Figure 13: Comparison of the quantity $n_{\textrm{eff}}$e($kR_{0}$) between the analytic spectrum obtained in Sec. V.4 using $n_{\textrm{eff}}=1$ (empty markers) and the numerical spectrum computed in Sec. VI by 3D FDTD using the usual effective index formula, for instance Eq. (4) of [28] (filled markers, TE modes in black and TM modes in red).
@@ VII Conclusion
In summary, the tractoid is integrable because of its rotational symmetry and in spite of the exponential divergence of ray trajectories induced by the negative curvature. When breaking the rotational symmetry, however, chaotic dynamics is expected [7], which will be the focus of a further exploration in non-Euclidean photonics.
@@ Acknowledgements.
This work was done with the C2N micro nanotechnologies platform. It was supported by the French RENATECH network, by the departement de l’Essonne, and by the French state funds managed by the ANR within the Investissements d’Avenir programme under reference ANR-19-CE09-0017 FRONTAL. S.B. acknowledges support for the Chair in Photonics from Région Grand-Est, Département de la Moselle, European Regional Development Fund, CentraleSupélec, Fondation CentraleSupélec, and the Eurometropole de Metz. B.D. acknowledges financial support from the Institute for Basic Science in Korea through the project IBS-R024-D1.
@@
J.-F. Audibert is acknowledged for the fabrication of centimeter-scale tractoids with a 3D printer, Y.-H. Melka and O. Ivashtenko for preliminary works on spherical squares, A. Comtet for Sec. V.2, and S. Nonnenmacher for Fig. 13.
Appendix A Bulk, effective, and group indices
This appendix discusses the different refractive indices relevant for the microlasers presented here.
  • The bulk index is the refractive index of the material itself. The bulk index of the resist IP-G was measured 1.51 at a wavelength of 600 nm in Ref. [22]. It is probably slightly different in our experiments due to the addition of the laser dyes and to the fabrication process. But as this index is difficult to measure precisely and its variation does not change significantly the results, we used the value 1.51 for the 3D simulations of Sec. VI.
  • The effective index is the refractive index which is experienced by the light propagating in a guiding layer. This is why it is involved in Eq. (18). The effective index is lower than the bulk index, since it considers the part of the wave which propagates outside the guiding layer (that is in air with $n=1$ in our case). The effective index can be calculated depending on the geometry of the device, on the bulk indexes of the involved layers, and on the polarization. In App. V.A of Ref. [13], we showed that the curvature of the propagating layer does not modify the effective index for the range of parameters of our experiments. For instance, the effective index is 1.5 at 600 nm for TE modes and a thickness of 1 $\mu$m for the most confined modes.
  • The group index takes into account the dispersion and must then be involved in formula (12) which considers the spacing between different wavelengths. A demonstration is given in [29] for instance. A calibration protocol is detailed in App. V.B of Ref. [13] and yields $n_{g}=1.58\pm 0.05$ for our experiments.